what is the magnetic induction at appoint 6m from the north pole of strength 150 amp-meter


1
Expert's answer
2021-03-31T16:14:47-0400
"B=\frac{\mu_0}{4\pi}\frac{2M}{r^3}\\
B=\frac{4\pi\cdot10^{-7}}{4\pi}\frac{2(150)}{6^3}=1.4\cdot10^{-7}T"